OpenStudy (anonymous):

Which of the following is the explicit formula for the sequence below? 8, 11, 14, 17, ... A. an = 8 + (n - 1)3 B. an = 8 + 3n C. an = 3 + 8n D. an = 3 + (n - 1)8

OpenStudy (dumbcow):

\[a_{n} = a_{1} +d(n-1)\] a1 = 8 d = 11-8 = 3